Sam mixed 91 mL of a acidic solution with 7 mL of a 58% acidic solution, to make a 37% acidic solution. Find the percent acid concentration of the first solution.
Answer by ankor@dixie-net.com(22740)

Sam mixed 91 mL of a acidic solution with 7 mL of a 58% acidic solution, to make a 37% acidic solution.
Find the percent acid concentration of the first solution.
:
Let x = the percent (in decmals) of the 1st solution
91x + .58(7) = .37(98)
91x + 4.06 = 36.26
91x = 36.26 - 4.06
91x = 32.2
x = 32.2/91
x = .354 which is a 35.4% 1st solution